In a right triangle the lengths of the legs are 5 ft and 12 ft. Find the length of the hypotenuse.
Answer by stanbon(75887)

In a right triangle the lengths of the legs are 5 ft and 12 ft. Find the length of the hypotenuse.
---
Equation:
h^2 = 5^2 + 12^2
---
h^2 = 25 + 144
---
h^2 = 169
----
h = 139
============
Cheers,
Stan H.